(1 point) If A and B are 3 x 3 matrices, det(A) = -5, det(B) = 9, then det(AB) = det(-2A) = det(AT) = det(B-1) = det(B3) =
Answer #1

det (A) -5 JAl= det (8) |Bl = det(@). detls) -519 - -45 io det (AB) = bez xAl = x det (2A) = -2A) = A A is nxn matai x. - -8